# Dust in the wind of outbursting young stars

### v1

@@ 3.2 Outburst mechanism and dusty winds
In Fig. 2, we zoom into one particular burst occurring at approximately 0.3 Myr. The duration of the luminosity burst in panel a) is years, which matches very well with canonical estimates of observed FU Ori events in YSOs (Audard et al., 2014; Fischer et al., 2023). The second panel shows the total () as well as thermal () ionization fractions, as measured at the inner boundary of the computational domain, at 0.51 au. The total ionization fraction includes contribution from cosmic rays as well as short-lived radionuclides. As both of these quantities are measured at the inner boundary, they show a marginally longer duration of outburst, as compared to the luminosity. Note that the degree of thermal ionization starts rising before the luminosity outburst. In our current model, the in the disk increases to the outbursting value of 0.1 above a threshold value of . During the outburst, rises steeply to maximum values that are consistent with total ionization of sodium at solar abundance (Balduin et al., 2023), dominating the total ionization fraction . Thus, we find that the results are not sensitive to the exact value of the chosen threshold in the model. The third panel of this figure shows parameter at the inner boundary. It starts rising slowly before the outburst, since the ionization balance equation results in a progressively larger active layer thickness. As the threshold value of is exceeded, achieves the maximum outbursting value of 0.1.
@@
Panel d) of Fig. 2 shows the mass accretion rate onto the star (), as well as mass ejection rates in both dust and gas due to disk winds, integrated over the centrifugal disk. It can again be seen clearly, that the wind gas mass loss rate is roughly equal to the accretion rate onto the star at quiescence, while during the outburst, this rate is about 10 times smaller. The disk ejects more dust during an outburst as compared to that in quiescence. This can be observed in the fourth panel, where the ratio of dust to gas mass loss rate in the wind () is a proxy for the wind dust-to-gas ratio. During quiescence, the winds carry much less than the canonical ISM value (dust-to-gas ratio 0.01) of dust, as most of the dust mass resides as grown dust. However, during an outburst, increases by over an order of magnitude, since a significant fraction of the grown dust is converted into small dust (see last panel of Fig. 1). Thus, our model predicts significantly more dusty winds during an outburst, which should eventually show as an extended dust emission beyond the disk. The increased production of small dust during an outburst can affect the ionization degree and in particular, the gas column density of the MRI-active region depends strongly on the ionization fraction of electrons. This is taken into account in the ionization balance equation by considering the rate of recombination onto the two dust populations of dust grains (Vorobyov et al., 2020b). The last panel e) of Fig. 2, implies that quiescent winds are extremely dust-depleted, which is consistent with our understanding of disk outflows (Suzuki et al., 2010; Gorti et al., 2015). Although the disk winds during outbursts are ”dusty” as compared to those in quiescence, they are still dust-depleted as compared to canonical ISM value. The total amount of small dust ejected during a single event is typically of the order of . This is a very small fraction of the dust reservoir available for planet formation, as the parent cloud core of mass 0.83 contains 1% of dust.
@@
Fig. 3, shows the progression of the outburst in terms of some key azimuthally averaged quantities. Here we aim to characterize the burst mechanism as well as the dust in the winds during an outburst, by showing the structural evolution. The three instances of time shown here correspond to the three vertical dashed lines plotted in Fig. 2, which represent the disk before, during (near the luminosity peak) and after an outburst. The first three panels depict the change in the disk structure across the outburst. Panel a) shows the gas surface density, which shows that the gaseous ring at 1.5 au is fully accreted onto the central star. Panel b) shows destruction of grown dust during the outburst, as most of the dust gathered in the pressure maximum at 1.5 au in the outbursting region shows flattening of the surface density profile. As most of the dust mass in a given cell is contained within the grown dust, its destruction leads to a significant increase in the small dust component. We see this in the panel c), where the small dust surface density increases by over two orders of magnitude in the outbursting region, which extends approximately 3 au.
@@
The next four panels of Fig. 3 focus on the mechanism of the outburst and production of small dust. The MRI-type outbursts occur naturally in our disk simulations and are already studied in detail (Kadam et al., 2020; Vorobyov et al., 2020b). In summary, the small viscosity in the dead zone is sufficient to increase the temperature in the innermost region. When it reaches ionization temperature of alkali metals, the ionization fraction increases exponentially and the affected region becomes prone to MRI instability, thus triggering an outburst. The increase in the midplane temperature during an outburst is seen in Panel d), while resulting increase in the thermal ionization fraction is seen in Panel e). The resulting increase in the disk to the maximum outbursting value of 0.1 is seen in Panel f). Thus, we show that the fundamental mechanism of the outburst remains unchanged with our current model of as a function of ionization fraction. Both the increase in temperature and during the outburst contribute towards a significant decrease in the fragmentation size of the dust particles. In the inner disk, the dust growth is fragmentation-limited and thus, controls the maximum size of the grown dust component (). As seen in panel g), decreases nearly three orders of magnitude during the burst in the inner disk. This essentially causes the aforementioned destruction of the grown dust component. The small dust is thus suddenly available for in large quantity in the outbursting region of the inner disk.
@@
The next four panels of Fig. 3 focus on the properties of the ejected winds. The panels h) and i) show the surface density loss rate due to winds for the gas and dust, respectively. Near the peak of the outburst, increases throughout the disk, as per its dependence on the luminosity of the central star. The dust loss rate shows a similar increase in the outer disk, however, in the outbursting region increases significantly. This is a result of the sudden production of the small dust, which is available to be ejected via winds. Panels j) and k) show the cumulative mass loss rate (, where is the area element) due to gas and dust, respectively. As expected, both of these quantities increase significantly during the outburst. Now consider the vertical lines in these plots, which show the region enclosing 50% and 90% of the cumulative mass loss rate. For the gas in the disk wind, the region within which a certain percentage of the ejected gas mass is contained remains relatively unchanged. In case of the dust component, however, the mass loss rate during an outburst is highly centrally concentrated, with 90% of it coming from au of the inner disk. This reflects the ejection of large amount of dust mass from the affected innermost region during an outburst.
@@
Lastly, as seen in panel l), the wind velocities are of the order of 10 of , which are consistent with observations of LVC components of disk winds (Natta et al., 2014; McGinnis et al., 2018; Fang et al., 2023). The outwards decreasing profile can possibly produce nested onion-like morphology of resolved sources (Pascucci et al., 2023). We make a rough estimate of the wind velocities by equating the work done by the wind torque to the kinetic energy gained by the mass ejected from the disk (see appendix A of Kadam et al. (2025)). During the outburst, the wind velocity does not increase, but rather decreases by approximately a factor of two. This is because the enhanced wind carries a relatively large amount of mass as compared to wind torques, which results in a lower kinetic energy per unit mass, in turn lowering the escape velocity. This decrease is analogous to the anti-correlation between the ejected mass and the associated velocity of the winds during coronal mass ejection events (Shibata & Magara, 2011). Although the wind velocity is estimated in a naive fashion, e.g., without taking into account thermal effects or wind interactions, the available observations on outflow forces also support the idea of very low wind velocities during FUor outbursts (Cruz-Sáenz de Miera et al., 2023).
@@ 4 Conclusions and discussions
In this paper, we presented results from our numerical experiments of modelling the formation and evolution of PPDs, with a focus on episodic accretion and ejection of dust during such outbursts via magnetic winds. The viscous in the inner disk during the outbursts is modelled as a function of the thermal ionization fraction. The disk winds are included with the formulation of Kadam et al. (2025), wherein during the outbursts, both the wind mass loss rate and torque increase as a function of accretion luminosity. We showed the mechanism of the outbursts in detail, which is identical to MRI-type bursts and the burst properties remain unaltered across evolutionary time. The qualitative agreement between an individual MRI-type eruption against observed longer duration FU Ori outbursts is excellent. This includes a fast rise time of yr, slow decline with a total duration of 100 yr, luminosity of few hundred and peak accretion rate of
@@
The MRI-type outbursts arise from residual viscous heating in the dead zone, which gradually results in midplane temperatures high enough to thermally ionize alkali metals. This leads to a sudden increase in the ionization fraction and thus effective , as the disk becomes MRI-active. During the outburst, the increase of temperature and in the affected inner disk region result in a lowering of the dust fragmentation size, and hence, destruction of the grown dust. The additional small grains suddenly become available for ejection with the magnetic winds, as only the small particles are aerodynamically coupled and entrained in the winds. One caveat of the model is that dust is considered to be formed of icy aggregates (small silicate monomers held together by water ice), which are destroyed instantaneously with an increase in temperature. However, such dust particles may be resilient to destruction at the onset of accretion events (Houge et al., 2024), hence we will consider a “slow” collisional fragmentation model in the future, to confirm the robustness of our findings. Another limitation of our simulations is that they are performed in two dimensions with thin-disk approximation and we have no information of the winds once they leave the computational domain. The extension of our model into the third dimension to track the disk winds remains a task for future exploration.
@@
During an outburst, the dust mass loss rate integrated over the disk increases by almost two orders of magnitude, as compared to the quiescent state. However, the winds during outburst are still about an order of magnitude dust-depleted as compared to the ISM. The innermost regions affected by the outburst contribute most to the dust mass loss rate, with 90% of the dust coming from innermost few au. The quantity of dust ejected does not significantly affect the mass reservoir available for planet formation. Since direct observations of outflows tracing dust emissions during FUor outbursts are rare, a concrete comparison with observations is an objective to be addressed in future work. However, our model of disk winds and resulting outflows during eruptions is in alignment with several observations associated with YSOs and young eruptive stars. For example, wind mass loss rate matching central accretion rate during quiescence, wind mass loss rate approximately 10% during eruption, dust-depleted winds during quiescence, entrainment of sub-micron-sized dust by the winds, decreased velocity of outflows during eruptions, the possibility of nested onion-like morphology with outwards decreasing velocity profile, and finally, increased dust content in the winds during eruptions. Our study sheds a light on the mechanism of dust production during outbursts and makes an order of magnitude predication on the dust mass loss rates in outflows and its origin in the innermost regions. Dusty winds during outbursts should have characteristic signatures as a periodic increase in the dust continuum emissions of the system.
@@ Acknowledgements.
- Lesur et al. (2023) Lesur, G., Flock, M., Ercolano, B., et al. 2023, 534, 465

### v2

@@ 3.2 Outburst mechanism and dusty winds
In Fig. 2, we zoom onto one particular burst occurring at approximately 0.3 Myr. The duration of the luminosity burst in panel a is years, which is very well matched with canonical estimates of observed FU Ori events in YSOs (Audard et al., 2014; Fischer et al., 2023). The second panel shows the total () as well as thermal () ionization fractions, as measured at the inner boundary of the computational domain, at 0.51 au. The total ionization fraction includes contribution from cosmic rays as well as short-lived radionuclides. As both of these quantities are measured at the inner boundary, they show a marginally longer duration of outburst, as compared to the luminosity. We note that the degree of thermal ionization starts rising before the luminosity outburst. In our current model, the value of in the disk increases to the outbursting value of 0.1 above a threshold value of . During the outburst, rises steeply to maximum values that are consistent with total ionization of sodium at solar abundance (Balduin et al., 2023), dominating the total ionization fraction, . Thus, we find that the results are not sensitive to the exact value of the chosen threshold in the model. The third panel of this figure shows parameter at the inner boundary. It starts rising slowly before the outburst, since the ionization balance equation results in a progressively larger active layer thickness. Once the threshold value of is exceeded, achieves the maximum outbursting value of 0.1.
@@
Panel d of Fig. 2 shows the mass accretion rate onto the star (), as well as the mass ejection rates in both dust and gas due to disk winds, integrated over the centrifugal disk. Again, it can be clearly seen that the wind gas mass-loss rate is roughly equal to the accretion rate onto the star at quiescence, while during the outburst, this rate is about ten times smaller. The disk ejects more dust during an outburst as compared to that in quiescence. This can be observed in the fourth panel, where the ratio of dust to gas mass-loss rate in the wind () is a proxy for the wind dust-to-gas ratio. During quiescence, the winds carry much less than the canonical ISM value (dust-to-gas ratio 0.01) of dust, as most of the dust mass resides as grown dust. However, during an outburst, increases by over an order of magnitude, since a significant fraction of the grown dust is converted into small dust (see last panel of Fig. 1). Thus, our model predicts significantly more dusty winds during an outburst, which should eventually show as an extended dust emission beyond the disk. The increased production of small dust during an outburst can affect the ionization degree and in particular, the gas column density of the MRI-active region depends strongly on the ionization fraction of electrons. This is taken into account in the ionization balance equation by considering the rate of recombination onto the two dust populations of dust grains (Vorobyov et al., 2020b). The last panel (e) of Fig. 2 implies that quiescent winds are extremely dust-depleted, which is consistent with our understanding of disk outflows (Suzuki et al., 2010; Gorti et al., 2015). Although the disk winds during outbursts are ”dusty” compared to those in quiescence, they are still dust-depleted as compared to canonical ISM value. The total amount of small dust ejected during a single event is typically of the order of . This is a very small fraction of the dust reservoir available for planet formation, as the parent cloud core of mass 0.83 contains 1% of dust.
@@
Figure 3 shows the progression of the outburst in terms of some key azimuthally averaged quantities. Here, we aim to characterize the burst mechanism as well as the dust in the winds during an outburst by showing the structural evolution. The three instances of time shown here correspond to the three vertical dashed lines plotted in Fig. 2, which represent the disk before, during (near the luminosity peak), and after an outburst. The first three panels depict the change in the disk structure across the outburst. Panel a shows the gas surface density, which shows that the gaseous ring at 1.5 au is fully accreted onto the central star. Panel b shows the destruction of grown dust during the outburst, as most of the dust gathered in the pressure maximum at 1.5 au in the outbursting region shows flattening of the surface density profile. As most of the dust mass in a given cell is contained within the grown dust, its destruction leads to a significant increase in the small dust component. We see this in the panel c, where the small dust surface density increases by over two orders of magnitude in the outbursting region, which extends approximately 3 au.
@@
The next four panels of Fig. 3 focus on the mechanism of the outburst and production of small dust. The MRI-type outbursts occur naturally in our disk simulations and are already studied in detail (Kadam et al., 2020; Vorobyov et al., 2020b). In summary, the small viscosity in the dead zone is sufficient to increase the temperature in the innermost region. When it reaches the ionization temperature of alkali metals, the ionization fraction increases exponentially and the affected region becomes prone to MRI instability, thus triggering an outburst. The increase in the midplane temperature during an outburst is seen in Panel d, while resulting increase in the thermal ionization fraction is seen in Panel e. The resulting increase in the disk to the maximum outbursting value of 0.1 is seen in Panel f. Thus, we show that the fundamental mechanism of the outburst remains unchanged with our current model of as a function of ionization fraction. Both the increase in temperature and during the outburst contribute towards a significant decrease in the fragmentation size of the dust particles. In the inner disk, the dust growth is fragmentation-limited and thus, controls the maximum size of the grown dust component (). As seen in panel g), decreases by nearly three orders of magnitude during the burst in the inner disk. This essentially causes the aforementioned destruction of the grown dust component. The small dust is thus suddenly available for in large quantity in the outbursting region of the inner disk.
@@
The next four panels of Fig. 3 focus on the properties of the ejected winds. Panels h and i show the surface density loss rate due to winds for the gas and dust, respectively. Near the peak of the outburst, increases throughout the disk, as per its dependence on the luminosity of the central star. The dust loss rate shows a similar increase in the outer disk, however, in the outbursting region increases significantly. This is a result of the sudden production of the small dust, which is available to be ejected via winds. Panels j and k show the cumulative mass-loss rate (, where is the area element) due to gas and dust, respectively. As expected, both of these quantities increase significantly during the outburst. We go on to consider the vertical lines in these plots, which show the region enclosing 50% and 90% of the cumulative mass-loss rate. For the gas in the disk wind, the region within which a certain percentage of the ejected gas mass is contained remains relatively unchanged. In case of the dust component, however, the mass-loss rate during an outburst is highly centrally concentrated, with 90% of it coming from au of the inner disk. This reflects the ejection of large amount of dust mass from the affected innermost region during an outburst.
@@
Lastly, as seen in panel l, the wind velocities are of the order of 10 of , which are consistent with observations of LVC components of disk winds (Natta et al., 2014; McGinnis et al., 2018; Fang et al., 2023). The outwards decreasing profile can possibly produce nested onion-like morphology of resolved sources (Pascucci et al., 2023). We make a rough estimate of the wind velocities by equating the work done by the wind torque to the kinetic energy gained by the mass ejected from the disk (see appendix A of Kadam et al. (2025)). During the outburst, the wind velocity does not increase, but rather decreases by approximately a factor of two. This is because the enhanced wind carries a relatively large amount of mass as compared to wind torques, which results in a lower kinetic energy per unit mass, in turn, lowering the escape velocity. This decrease is analogous to the anti-correlation between the ejected mass and the associated velocity of the winds during coronal mass ejection events (Shibata & Magara, 2011). Although the wind velocity is estimated in a naive fashion, for instance, without taking into account thermal effects or wind interactions, the available observations on outflow forces also support the idea of very low wind velocities during FUor outbursts (Cruz-Sáenz de Miera et al., 2023).
@@ 4 Conclusions and discussions
In this paper, we presented results from our numerical experiments of modeling the formation and evolution of PPDs, with a focus on episodic accretion and ejection of dust during such outbursts via magnetic winds. The viscous in the inner disk during the outbursts is modeled as a function of the thermal ionization fraction. The disk winds are included with the formulation of Kadam et al. (2025), wherein during the outbursts, both the wind mass-loss rate and torque increase as a function of the accretion luminosity. We have shown the mechanism of the outbursts in detail, which is identical to MRI-type bursts and the burst properties remain unaltered across evolutionary time. The qualitative agreement between an individual MRI-type eruption against observed longer duration FU Ori outbursts is excellent. This includes a fast rise time of yr, slow decline with a total duration of 100 yr, luminosity of few hundred , and peak accretion rate of
@@
The MRI-type outbursts arise from residual viscous heating in the dead zone, which gradually results in midplane temperatures high enough to thermally ionize alkali metals. This leads to a sudden increase in the ionization fraction and thus effective , as the disk becomes MRI-active. During the outburst, the increase of temperature and in the affected inner disk region result in a lowering of the dust fragmentation size and, hence, the destruction of the grown dust. The additional small grains suddenly become available for ejection with the magnetic winds, as only the small particles are aerodynamically coupled and entrained in the winds. One caveat of the model is that dust is considered to be formed of icy aggregates (small silicate monomers held together by water ice), which are destroyed instantaneously with an increase in temperature. However, such dust particles may be resilient to destruction at the onset of accretion events (Houge et al., 2024); hence, we plan to consider a “slow” collisional fragmentation model in the future to confirm the robustness of our findings. Another limitation of our simulations is that they are performed in two dimensions with thin-disk approximation and we have no information of the winds once they leave the computational domain. The extension of our model into the third dimension to track the disk winds remains a task for future exploration.
@@
During an outburst, the dust mass-loss rate integrated over the disk increases by almost two orders of magnitude, as compared to the quiescent state. However, the winds during outburst are still about an order of magnitude dust-depleted as compared to the ISM. The innermost regions affected by the outburst contribute most to the dust mass-loss rate, with 90% of the dust coming from innermost few au. The quantity of dust ejected does not significantly affect the mass reservoir available for planet formation. Since direct observations of outflows tracing dust emissions during FUor outbursts are rare, a concrete comparison with observations is an objective to be addressed in future work. However, our model of disk winds and resulting outflows during eruptions is in alignment with several observations associated with YSOs and young eruptive stars. For example, a wind mass-loss rate matching the central accretion rate during quiescence, wind mass-loss rate of approximately 10% during eruption, dust-depleted winds during quiescence, entrainment of sub-micron-sized dust by the winds, decreased velocity of outflows during eruptions, the possibility of nested onion-like morphology with outwards decreasing velocity profile, and, finally, increased dust content in the winds during eruptions. Our study sheds light on the mechanism of dust production during outbursts and makes a predication of an order of magnitude on the dust mass loss rates in outflows and its origin in the innermost regions. Dusty winds during outbursts are expected to exhibit characteristic signatures, such as a periodic increase in the dust continuum emission of the system.
@@ Acknowledgements.
- Lesur et al. (2023) Lesur, G., Flock, M., Ercolano, B., et al. 2023, in Astronomical Society of the Pacific Conference Series, Vol. 534, Protostars and Planets VII, ed. S. Inutsuka, Y. Aikawa, T. Muto, K. Tomida, & M. Tamura, 465